Resolving power?

Grade:11

1 Answers

Shyam Chand
10 years ago
the resolution (resolving power) of a microscope means its ability to distinguish two items at its highest magnification. the same goes for any other optical instrument. its like watching twolines which are extremely close to each other with unaided eye and then watching them with the microscope with the unaided eye, they will appear as one line with the microscope they will appear distinct. If d? is the angular separation, then Resolving Power is given by the formula R= 1/d? = D/1.22? ; whereDis the aperture of the objective; ?is the wavelength of the light .